OpenStudy (anonymous):

can any one help me in H8...????

OpenStudy (anonymous):

H8P2..??

OpenStudy (anonymous):

all.............

OpenStudy (nurali):

@vivekgarg 1) 687 2) .193 3) 643 4) .5011 5) .6 6) 3.45 7) 1.8

OpenStudy (nurali):

@Neelima 1. 5.7143 2. 296.783 3. 8 4. 166.667

OpenStudy (nurali):

@Neelima if correct please click best response

OpenStudy (anonymous):

@Nurali what is the formulae for last two qns of H8P3

OpenStudy (anonymous):

@Neelima: H8p1: 1) 0 2)2000 3)270.6706 4)3.0303e3 5)146.3667 I got this answer through Matlab.......

OpenStudy (anonymous):

@Neelima: h8p2: 1) 5.7344 2) 296.78 3) 8 4) 166.66

OpenStudy (anonymous):

H8P3 i stuck on the fourth answer....the others are a)185.347405561 b)4.99926326647 c)3.99926326647 d)???????? e)29.5159549597

OpenStudy (anonymous):

wats d formula for the last one??

OpenStudy (anonymous):

hey and body plz formulae for last twoqns of H8P3

OpenStudy (anonymous):

Q4 VD1 : high; VG2=VOL; Then VStore: high à Q3:on; How long does it take to change VG2 to reach VOH? Again we can use the generic dynamics of a 1st order system: VG2(t) = VG2(inf) + (VG2(0) - VG2(inf)) * exp(-t/τ) VOH = VG2max + (VOL - VG2max) * exp(-t/τ) 3.5 = 4 + (1.5 – 4) * exp(-t/τ) t = 1.61 * τ Regarding the value of time constant τ, the resistance that Cgs sees should be calculated. In this case since Q1 is off and Q3 is on, R = RON3 + RPU || ROFF1 τ = RC t= 1.61 * τ and then i get wrong result... wtf...

OpenStudy (anonymous):

can anyone give the formulae for H8P3

OpenStudy (anonymous):

H8p3 last 2 qns formulae ;D 3RD---(-ln( (3.5-5) / (0.5-5) ) * (RPU+RON) * 3.5*10^-15) * 10^12 4TH----(-ln( (0.9-(5*1800/(RON+RPU))) / (3.5-(5*RON/(RON+RPU))) ) * ((RON|| RPU) + RON) * 3.5*10^-15) * 10^12